OpenStudy (anonymous):

find the linear approximation of sqrt(8+x) where a=8 and delta x=-1

OpenStudy (freckles):

The tangent line to the function f(x)=sqrt(8+x) at x=a is an approximation to the function f(x)=sqrt(8+x) for values near x=a.